1)if a+b=3-cos4 theta and a-b=4sin theta  then ab is always less than or equal to
 
options a) 1/2
b)1
c)2/3
d)3

Grade:10

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
Can you please confirm if you have:
a+b=3-cos\left ( 4\Theta \right ), a+b=4sin\left ( \Theta \right )
I feel there is some mistake in this.
Please correct the same.
Regards
Sumit